OpenStudy (anonymous):

Find two numbers whose product is 15 such that one of the numbers is 7 more then 4 time the other number.

OpenStudy (agreene):

xy=15 x+7=4y solve system

OpenStudy (anonymous):

x + y = 15 y = 15 - x 7 + 4y = x 7 + 4(15 - x) = x 67 - 4x = x 67- 67 - 3x = -67 -3x = -67 (divide both sides by -3) x = 22.33 y = 15 - 22.33 y = -7.33 i think.. im not too sure

OpenStudy (anonymous):

xy = 15 x = 4y + 7 substitute in (4y + 7)y = 15 4y^2 + 7y = 15 4y^2 + 7y - 15 = 0 (4y - 5)(y + 3) = 0 y = 5/4 or y = -3 numbers -3 and -5 number 12 and 1 1/4

OpenStudy (anonymous):

p yea this ^ is ryt i didn't see it's product my bad